Journalist Report – February 27th

Sol 15 – Cosmic Chasms

“The men on the porch listened. Hearing nothing, they extended their thoughts and their imaginations out and into the surrounding meadows.”

– Chapter 15 of The Martian Chronicles by Ray Bradbury

It’s back to normal for us today, if anything about this mission can be considered “normal”! Corentin got us up and running to start the day with stretching and leg exercises, and we enjoyed breakfast as a crew in the Upper Deck before Alice, Quentin and Corentin had to get ready for the first EVA of the week. The photogrammetry experiment continues! Today’s destination was Candor Chasma, a canyon not far from the station and a place the previous ISAE-Supaero crew had mentioned to us as the best destination for an exploration EVA. The sinuous paths and harsh landscape gave the three of them the opportunity to choose difficult checkpoints to find, which we hope will amplify the added value of the 3D map, compared to its 2D equivalent. After choosing 10 checkpoints, the EVA crew had time to explore Candor Chasma. Thankfully for them, the strong winds we experienced last night dwindled in the early morning, and got stronger after they returned to the Hab, allowing them to use the Parrot drone to 3D-map the canyon. As they ventured out further and further away from the station, Adrien serving as HabCom gradually lost communication with them, and could only “extend his thoughts and imagination” to Candor Chasma with the help of a few bits of radio chatter caught here and there.
While the crew was out on EVA, I visited Alexandre in the observatory. Because of the weather conditions in the past two weeks, very little use of the two MDRS telescopes observatories was made. As the sky gets clearer, Alexandre will be able to start his astronomy project. His objective is to characterize little known asteroids to contribute to an American database. By analyzing the light curve of an asteroid (meaning its luminosity as a function of time), you can deduce some of its characteristics, such as its rotation period, its size, etc. Alexandre’s first observations focused on asteroid Eos. Nevertheless, he was frustrated by the unrelenting wind, as the conditions prevented him from opening the dome of the Musk Observatory and observing the Sun. The observations he had programmed last night were also not performed by the robotic observatory…
In the afternoon, Quentin deployed all the location tracking anchors in the different modules. I toured the station with him as I wore my tag, and he measured the distance between me and the different anchors. Back at the Hab, we checked the data recorded by the trackers, and found that our movements were correlated with it! Quentin is glad to announce that the location tracking system is now fully deployed and operational.
Adrien and Corentin also conducted a large-scale harvest in the GreenHab: dill, spinach and other greens will make our dinner a little more earthly…

Astronomy Report – February 27th

Astronomy Report

Name: Alexandre VINAS

Crew: 275
Date: 02/27/2023


Robotic Telescope Requested (choose one) MDRS-14 MDRS-WF

Objects Viewed: /

Problems Encountered: /


Solar Features Observed: /

Problems Encountered: The dome has not been open because of the high winds. I only made a « solar alignment », and tested the camera and the equipment, in order to prepare the next solar observation day.

EVA Report – February 27th

Crew 275 EVA Report 27Feb2023

EVA # 13

Author: Alice Chapiron, EVA Leader and Crew Scientist

Purpose of EVA: The purpose of the EVA was to perform photogrammetry at Candor Chasma using drones.

Start time: 9:05

End time: 12:49


09:05: Airlock closed.

09:11: Rovers Opportunity and Perseverance unplugged. We then left from the Hab at 09:14.

09:16: Rovers parked near the atmospheric instruments.

09:55: Data retrieved, and batteries changed for the Purple Air, LOAC, Mega-Ares, and Weather Station. The EVA crew is about to go back to the rovers to go to their target location (Candor Chasma)

10:03: Rovers are parked at the entrance of Gateway to Candor, the EVA Leader switches to channel 2 to communicate properly with her EVA crewmates, then the EVA Crew begins to walk towards their destination.

10:25: The EVA Crew enters Candor Chasma

10:45: The EVA Crew arrived at destination (further point they will go in Candor Chasma). The wind allowing us to fly the drone in Candor Chasma, the EVA crew walks back doing photogrammetry of Candor Chasma and defining the checkpoints.

12:10: The EVA Crew is tidying up their material. They are walking back to the rovers.

12:35: EVA Crew back to the Rovers.

12:41: Arrival at the Hab.

12:49: End of the EVA.

Destination: Candor Chasma

Coordinates: N4251000 E521500

Participants: EVA Leader: Alice Chapiron (Crew Scientist), Quentin Royer (Crew Engineer), and Corentin Senaux (HSO)

HAB COM: Adrien Tison (Crew Botanist)

Road(s) and routes per MDRS Map: Cow Dung Road until the intersection with “Gateway to Candor”, we then walked eastwards until Candor Chasma for an estimated 2.5 km from the rovers.

Mode of travel: Driving using Perseverance and Opportunity and walking

GreenHab Report – February 27th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 75.4°F

Average humidity: 23 %

Hours of supplemental light: 0 hour

Daily water usage for crops: 12.55 gallons

Daily water usage for research and/or other purposes: 0.13gallons

Water in the blue tank: 161.94 gallons

Time(s) of watering for crops: 9:23 & 14:45

Changes to crops: I followed the instructions given by Shannon and took care of the dill and greens. I then cut the dills and pulled half of it. I harvested the greens in the planters and replanted new seeds.

Narrative: Today, I had more workload than usual but for the better! This morning I was the HabCom for today’s EVA so it was the crew astronomer and crew journalist who watered the plants and crops in the morning.

But many things were to be done, so I went back to the GreenHab in the afternoon. Then, I began by watering plants and crops to ensure their well-being. And finally, I started harvesting greens, lettuce, kale, and dills with the help of the HSO.

We took care of the dills and after a long work we finished and replanted new greens seeds.

Harvest: 102 g of kale, 254 g of greens, 30 g of lettuce, 124 g of dill

Support/supplies needed: N/A

Operations Report – February 27th

Crew 275 Operations Report 27Feb2023
SOL: 15

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• GreenHab Carbon Monoxide alarm

Notes on non-nominal systems: None

Spirit rover used: No
• Hours: 209.6
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: Yes
• Hours: 113.5
• Beginning charge: 100%
• Ending charge: 77%
• Currently charging: Yes

Curiosity rover used: No
• Hours: 221.4
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Perseverance rover used: Yes
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: 68%
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. Location tracking sensors were placed.
• WATER USE: 20.60 – The formula is inaccurate for the top of the tank
• Water: 457.82 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: No

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS. MDRScampus is used for experimental purpose.

Summary of suits and radios: Suits and radios were used during today’s EVA. Nothing to report.

Summary of GreenHab operations: The Crew botanist conducted routine operations. He harvested and reseeded greens. A location tracking sensor was placed.
• WATER USE: 12.68 gallons
• Heater: On
• Supplemental light: No
• Harvest: 102 g of kale – 254 g of microgreens – 30 g of lettuce – 124 g of dill

Summary of ScienceDome operations: No operations were conducted in the ScienceDome. Location tracking sensors were placed.
• Dual split: Switched off during the night

Summary of RAM operations: No operations were conducted in the RAM today. A location tracking sensor was placed.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns, and requests to Mission Support: None.

Sol Summary Report – February 27th

Crew 275 Sol Summary Report 27Feb2023

Sol: 15

Summary Title: Cosmic Chasma

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

Today we performed our very first 4-hour EVA at Candor Chasma. The aim of this ambitious EVA was to define checkpoints and perform photogrammetry of the area using Anafi drones. This time, the area of interest was much wider and required more pictures (the computer is still working on building the 3D render of the terrain). The EVA crew was amazed by the landscapes that they had the opportunity to see, and the four remaining crewmembers are looking forward to seeing that for themselves later during the week.

With the clear sky of last night, our crew astronomer thought that the observations he had planned would have been performed, but he was disappointed to see that it was not the case. Unfortunately, it was also too windy to observe the Sun today…

At the GreenHab, it was time for an intense harvest. New seeds have also been planted. We are all looking forward to seeing them grow as we get closer to the end of our mission every day.

Look Ahead Plan:

After today’s EVA, some post-processing is needed to prepare the two next EVAs to Candor Chasma. The first one would take place on Wednesday (Sol 17) and the second one on Friday (Sol 19). The EVA crews would have to find specific checkpoints based on a 2D map or a 3D render. We are also planning to have a second geology EVA on Sol 20, to further test the MetMet instrument. We would like to take the opportunity of these different EVAs to change the batteries of the weather instruments, since they cannot last for more than two days, especially in cold weather.

The BBC is planned to arrive at the MDRS on Friday (Sol 19) between 8:30 and 9:00. They would follow us on our EVA to Candor Chasma, then film inside the station in the afternoon and leave at around 17:00.

Anomalies in work: None.

Weather: Sunny and windy, low -2°C/28°F, high 8°C/46°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site